OpenStudy (anonymous):

MEDAL + FAN FOR RIGHT ANSWER A large container holds 5 gallons of water. It begins leaking at a constant rate. After 10 minutes, the container has 3 gallons of water left.

OpenStudy (anonymous):

A. at what rate is the water leaking? B. after how many minutes will the container be empty

OpenStudy (anonymous):

can u guys plz help

OpenStudy (anonymous):

2

OpenStudy (anonymous):

wait but do i divide 5 with 10 or 3

OpenStudy (anonymous):

1/5

OpenStudy (anonymous):

wait why did u delete it

OpenStudy (anonymous):

Initially at t=0 suppose is the instant when the container had 5 gallons of water. After time t=10mins, the container has 3 gallons, so water leaked =5-3 =2 gallons in 10 minutes hence rate of leak=2/10=1/5 gallons per minute

OpenStudy (anonymous):

okay wht about my second part

OpenStudy (anonymous):

i got 25 min

OpenStudy (anonymous):

rate of leak=1/5 gallons per minute. You need to find the time when leak=5 gallons. Now 1minute =1/5 gallons of leak hence 1gallon of leak==5 minute therefore 5 gallon of leak=5*5=25 minutes

OpenStudy (anonymous):

okay thts right can u help me with one more plzzz

OpenStudy (anonymous):

I will try, you tag me

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

A line passes through (6, 3), (8, 4), and (n, -2). Find the value of n? @Princer_Jones

OpenStudy (anonymous):

Ok first we need to construct the equation of the line from the first two points. Did you construct it? @hhmh4

OpenStudy (anonymous):

hold up

OpenStudy (anonymous):

-4 is the answer,

OpenStudy (anonymous):

okay thank you soooo much

OpenStudy (anonymous):

see the equation of the line connecting the first two points is (x-6)/(8-6)=(y-3)/(4-3) this implies x=2y. Putting y=-2 we have n=-4

OpenStudy (anonymous):

ohh yes